OpenStudy (anonymous):

3x-(-x+1)=[3(x+6-10]-[5-2(x-8)]... I've solved and stuck on... 4x-1=[3x+12]-[5-2x+16)] @Compassionate, @some_someone

OpenStudy (compassionate):

I'm eating right now so you may not get a response or it'll be slow.

OpenStudy (compassionate):

Wait about 10 minutes. Okay? If no one has helped you by then, then I will.

OpenStudy (compassionate):

@some_someone , help her while I'm eating. That's an order.

OpenStudy (anonymous):

yup. whatever works. as long as I get some help i'll be waiting here all night

OpenStudy (anonymous):

You should have 3x-12, not 3x+12

OpenStudy (anonymous):

\[4x-1=(3x-12)-(5-2x+16) \]

OpenStudy (anonymous):

okay . let me where that happens... 3x-(-x+1)=[3(x+6-10]-[5-2(x-8)] 3x+x-1=3x+18-30]-[3(x-8)]

OpenStudy (anonymous):

4x-1=[3x-12]-[3x-24]...okay I see it cool thanks!

OpenStudy (anonymous):

Do you now see how to solve it? If not I can show that too

OpenStudy (anonymous):

that would be so awesome and great.i can never get past [ ] things and doing these same problems but with fractions

OpenStudy (anonymous):

\[4x-1=5x-17-16\]

OpenStudy (anonymous):

\[4x-1=5x-33\]

OpenStudy (anonymous):

\[4x+32=5x\]

OpenStudy (anonymous):

\[32=x\]

OpenStudy (anonymous):

how did you get 5x-17-16??

OpenStudy (anonymous):

\[(3x-12)-(5-2x+16)\]

OpenStudy (anonymous):

\[3x-12-5+2x-16\]

OpenStudy (anonymous):

okay.got it

OpenStudy (anonymous):

Then you just kind of have to rearrange... and get \[3x+2x-17-16=5x-33\]

OpenStudy (anonymous):

Glad I could help! I was just trying to kill time waiting for someone to answer my chemistry question but I give up...

OpenStudy (anonymous):

oh well I do greatly appreciate your help.i wish I could help but im not that strong in math as you can see.